# Rethinking the effects of data contamination in Code Intelligence

### v2

@@
Abstract
In recent years, code intelligence has gained increasing importance in the field of automated software engineering. Meanwhile, the widespread adoption of Pretrained Language Models (PLMs) and Large Language Models (LLMs) has raised concerns regarding data contamination and its potential impact on model performance evaluation. This paper presents a systematic empirical study to investigate the fine-grained data contamination on code intelligence tasks. Our study involves diverse representative PLMs, namely RoBERTa and GPT-2, and LLMs, namely LLaMA and StarCoder, covering three major tasks: code translation, code generation, and code summarization. We categorize contamination scenarios into four types according to the code intelligence practice, namely input-only, output-only, unpaired, and paired contamination settings, and construct corresponding experimental and control groups for exploration.
@@ Abstract
Index Terms:
data contamination, code intelligence tasks, large language model.I Introduction
In recent years, code intelligence has been a critical research area in the software engineering field, significantly improving software development and maintenance automation [1, 2, 3, 4, 5, 6]. This is mainly attributed to the advancement of Pre-trained Language Models (PLMs) [7, 8, 9, 10, 11, 12, 13, 14] and Large Language Models (LLMs) [15, 16, 17, 18, 19, 20] on code, especially their effective learning on massive Natural Language (NL) and Programming Language (PL) corpora. Despite their success, researchers began to question whether PLMs or LLMs’ high performances stem from their memorization of pre-training data, i.e., data contamination, and many related studies were proposed [21, 22, 23, 24, 25]. Nevertheless, most of them concentrate on sample-level contamination, which means both input and output portions are leaked in pairs in PLMs/LLMs’ pre-training corpora.
@@
However, contamination does not always show up in samples, especially in the code intelligence field. For example, when doing code translation from Java to ArkTs, a declarative PL for application development in the HarmonyOS ecosystem [26]. PLMs/LLMs may not be pre-trained on ArkTS, but have seen massive Java code, including the one to be translated. In this case, a contamination scenario concerning only input portions occurs. Other scenarios are also prevalent in code tasks, such as the unit test case generation task. Perhaps both the focal methods and test cases have been pre-trained by PLMs/LLMs, but they are placed in different files on GitHub, leading to their unpaired learning during the pre-training stage of PLMs/LLMs. In this case, how does it affect PLMs/LLMs’ performance on downstream tasks? As such, this work seeks to fill this important gap.
@@
Empirical Study: To explore the contamination effects in a fine-grained manner, we conduct a comprehensive empirical study involving two PLMs and two LLMs on three representative code intelligence tasks: code translation, code generation, and code summarization. In our analysis, we identify four typical types of data contamination scenarios commonly encountered in code intelligence: (1) input-only contamination, where only the input portion of the test sets appears in the pre-training corpora; (2) output-only contamination, where only the output portion is contaminated; (3) unpaired contamination, where both input and output components are seen in pre-training but not as matching pairs; and (4) paired contamination, where the exact input-output pairs from the test sets are found in the pre-training data of PLMs or LLMs.
@@
As for experiments with PLMs, we adopt a pre-training, fine-tuning, and inference paradigm to simulate their practical usage. To circumvent uncontrolled data leakage, we perform pre-training from scratch [24, 27] and simulate contamination by injecting test data into the pre-training corpus according to different contamination settings. After fine-tuning on downstream tasks, we evaluate model performance on test sets and compare with the uncontaminated counterparts, thereby measuring the contamination effects. As for LLMs, we still follow their practical usage paradigm, i.e., pre-training then inference, to conduct the experiments. We use their publicly available pre-training corpora [16, 18] to extract appropriate samples according to different contamination settings as contaminated testing sets. As for uncontaminated counterparts, we design a series of perturbation rules to modify the contaminated testing sets, making them unseen to LLMs [28] while keeping sample complexities the same. All experiments are repeated five times for statistical analysis, thereby making comparisons and discussing the effects of contamination.
@@
Main Findings: (1) PLMs with pre-training, then fine-tuning and inference paradigm are surprisingly not significantly affected by any contamination settings, overall yielding on average 0.16% and -0.06% of fluctuations on BLEU and METEOR over uncontaminated counterparts (p-value0.05). (2) Direct inference or small-scale fine-tuning uncovers the contamination effects on decoder-only PLMs, such as GPT-2, achieving 73.03% and 47.18% of overall lifting in terms of BLEU and METEOR (p-value0.05). (3) Large-scale fine-tuning tends to eliminate the contamination effects of decoder-only PLMs by overwhelmingly learning with underlying knowledge and task alignment specifications. (4) LLMs with pre-training and inference paradigm are only affected by paired contamination, obtaining consistent lifting of 10.87% and 5.67% on both metrics overall (p-value0.05). This paper makes the following contributions.
@@
(3) We conduct extensive experiments among diverse contamination settings, various PLMs and LLMs on multiple code intelligence tasks. A series of findings are revealed in this work, shedding light on future research and practice.
II Experimental Setup
@@
II-A Research Questions
This paper proposes the following four research questions according to different contamination settings.
II-A1 RQ1: Does Input-Only Contamination Lead to Performance Overestimation?
When constructing testing sets, only the input portion is included in the pre-training corpora of PLMs/LLMs, thereby simulating the input-only contamination. This is a prevalent scenario, the Java-to-ArkTS translation is a typical example as mentioned in Section I.
A prior study conducted by Jiang et al. [27] has claimed that the input-only and output-only contamination leads to performance overestimation for PLMs in Natural Language (NL) tasks, such as SQuAD and MMLU. However, NL tasks differ significantly from code intelligence tasks, and their experimental designs are flawed. First, the NL tasks carried out in Jiang et al.’s study only involve English text, leading to a relatively high similarity between their inputs and outputs compared with code intelligence tasks that always contain both NLs and code from different PLs. Second, the experiments of Jiang et al. only make comparisons in terms of average performance among only three trials without conducting statistically significant tests, making it hard to eliminate the influence of randomness and determine whether the outperformance of data contamination is consistent and prominent. Third, Jiang et al. only conducted their experiments on one PLM, i.e., GPT-2, without considering different architectures of PLMs and other LLMs. Therefore, the input-only contamination is highly worthwhile to re-evaluate in the code intelligence domain.
II-A2 RQ2: Does Output-Only Contamination Lead to Performance Overestimation?
@@
II-A3 RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
@@
This is a kind of complete contamination of testing sets, but the input and output portions are not paired in the pre-training corpora of PLMs/LLMs. Recall the example we mentioned in Section I, when doing unit test case generation task, both focal methods and test cases may exist in the pre-training corpora of PLMs/LLMs, but their source files are split in the wild. Thus, these code samples will be pre-trained in PLMs/LLMs in a split manner. Considering the prevalence of this setting, unpaired input-output contamination is also meaningful and critical for exploration.
II-A4 RQ4: Does Paired Contamination Lead to Performance Overestimation?
@@
As another kind of complete contamination, paired input-output contamination represents that both the input and output portions of testing sets have been seen by PLMs/LLMs, and they are concatenated in pairs during the pre-training stage. Many code intelligence tasks are very likely to involve this contamination, such as code generation, as Natural Language Descriptions (NLD) are always closely attached in front of corresponding code snippets, leading to their paired appearance in PLMs/LLMs’ pre-training. Theoretically, paired contamination is the most likely type to cause performance overestimation, and the LLM part has been investigated in [23, 22], but PLMs with pre-training, fine-tuning, and inference paradigm have not been explored.
II-B Models
II-B1 Pre-trained Language Models (PLMs):
@@
Among this category, an encoder-only model, namely RoBERTa [29], and a decoder-only model, namely GPT-2 [30], are selected for experiments, as they have covered the most representative architectures of PLMs. We select RoBERTa with 125M parameters and GPT-2 with 150M parameters for experiments according to the load of our computational resources. To facilitate our research, we adopt one typical pre-training task for each PLM during its pre-training stage. RoBERTa is pre-trained with the Masked Language Modeling (MLM) [31] task, while the Causal Language Modeling (CLM) [32] task is used for GPT-2. Contamination injections are performed in the pre-training stage. Then, PLMs go through fine-tuning and inference stages to observe whether contamination induces performance overestimation compared with the uncontaminated counterparts.
II-B2 Large Language Models (LLMs):
@@
As for this category, we select one general LLM, namely LLaMA [16], and one code LLM, namely StarCoder [18]. LLaMA is a family of multilingual LLMs with parameter sizes ranging from 7B to 65B. It is pre-trained on 1.4 trillion tokens of various open-source datasets, where 4.5% of the training base is code extracted from GitHub [33]. We select its 33B version for experiments. StarCoder is a Code LLM containing 15.5B parameters, trained on over 1 trillion code tokens. LLMs are experimented with in a pre-training then inference paradigm. Due to the massive scale and substantial training base of LLMs, retraining from scratch is impractical. We curate appropriate samples from their pre-training corpora for the contaminated testing set construction and experiments. The reasons we consider the above two LLMs are two-fold: First, they cover both general and code LLMs, making our exploration comprehensive. Second, they are two of the few open-source LLMs that also release their pre-training corpora, enabling us to construct contaminated testing sets and the uncontaminated counterparts from them.
II-C Code Tasks and Metrics
II-C1 Code Translation:
@@
This task involves converting code from one PL to another while keeping the functionality the same, which is critical in scenarios such as migrating legacy systems and adopting modern frameworks. In this work, we focus on code translation between Java and C# for experiments. For simplicity, we use JavaC# to denote Java-to-C# translation while using C#Java to denote the opposite.
II-C2 Code Generation:
@@
It is a process of automatically producing source code from a given NL requirement, aiming to reduce manual coding effort while ensuring correctness and efficiency. In this work, we adopt Java code generation for experiments, which is denoted as NLJava for simplicity.
II-C3 Code Summarization:
@@
This task focuses on generating a NL description for a given code snippet, which is crucial for improving code readability, maintenance, and collaboration among developers. Still, we carry out experiments with code summarization in Java programs, which is denoted as JavaNL for simplicity.
II-C4 Metrics:
@@
To automatically assess the quality of the outputs for the above code tasks, we adopt two widely used evaluation metrics, i.e., BLEU [34] and METEOR [35]. BLEU measures the precision of -gram overlap between the generated output and the reference, focusing on how much of the generated content is correct. In contrast, METEOR incorporates stemming, synonym matching, and alignment-based evaluation, placing more emphasis on recall—i.e., how comprehensively the generated output covers the information in the reference. By combining these two complementary metrics, we provide a more balanced and holistic assessment of generation quality from both precision and recall perspectives.
II-D Datasets
The following paragraphs present the datasets used in this paper for contamination construction, which has been partitioned into two parts (i.e., PLM and LLMs) for illustration.
II-D1 Dataset preparation for PLMs:
As for the pre-training corpus, we use the Java portion of the widely adopted CodeSearchNet dataset [36], which contains over 160,000 Java function snippets for mono-lingual pre-training. The downstream tasks contain code translation and code generation for the experiments of PLMs. To achieve this, we adopt the CodeTrans dataset [10], which includes 10,300 training samples, 500 validation samples, and 1,000 test samples for code translation between Java and C#. Besides, we use the NL-to-Java code generation dataset, namely CONCODE [37], for experiments, which contains 100,000 training samples and 2,000 validation and testing samples, respectively. Since we pre-train PLMs from scratch, we can use the testing sets of diverse code tasks to deliberately inject contaminated samples into pre-training corpora, also easily construct uncontaminated counterparts. Detailed experimental methods are elaborated in Section II-E.
II-D2 Dataset preparation for LLMs:

### v3

@@
Abstract.
In recent years, code intelligence has gained increasing importance in the field of automated software engineering. Meanwhile, the widespread adoption of Pretrained Language Models (PLMs) and Large Language Models (LLMs) has raised concerns regarding data contamination and its potential impact on model performance evaluation. Previous studies mainly focused on sample-level contamination, ignoring partial contamination scenarios that are pervasive in code intelligence. This paper fills this gap and presents a systematic empirical study to investigate the fine-grained data contamination on mainstream code tasks. Our study involves diverse representative PLMs: RoBERTa and GPT-2, and LLMs: LLaMA and StarCoder, covering three major tasks: code translation, code generation, and code summarization, across two Programming Languages (PLs): Java and Python. We categorize contamination scenarios into four types according to the code intelligence practice, namely input-only, output-only, unpaired, and paired contamination settings, and construct corresponding experimental and control groups for exploration.
@@ Abstract.
1. Introduction
@@
In recent years, code intelligence has been a critical research area in the software engineering field, significantly improving software development and maintenance automation (Sun et al., 2025; Zhang et al., 2024; Jiang et al., 2024a; Alon et al., 2018, 2019; Zhang et al., 2019b). This is mainly attributed to the advancement of Pre-trained Language Models (PLMs) (Feng et al., 2020; Guo et al., 2020; Svyatkovskiy et al., 2020; Lu et al., 2021; Ahmad et al., 2021; Wang et al., 2021, 2023b; Guo et al., 2022) and Large Language Models (LLMs) (Chen et al., 2021; Touvron et al., 2023; Nijkamp et al., 2022; Li et al., 2023; Roziere et al., 2023; Lozhkov et al., 2024) on code, especially their effective learning on massive Natural Language (NL) and Programming Language (PL) corpora. Despite their success, researchers began to question whether PLMs or LLMs’ high performances stem from their memorization of pre-training data, i.e., data contamination, and many related studies were proposed (Cao et al., 2024; Ramos et al., 2024; Riddell et al., 2024; Magar and Schwartz, 2022; Golchin and Surdeanu, 2024). Nevertheless, most of them concentrate on sample-level contamination, which means both input and output portions are leaked in pairs in PLMs/LLMs’ pre-training corpora. However, contamination does not always show up in samples, especially in the code intelligence field. For example, when doing code translation from Java to ArkTs, a declarative PL for application development in the HarmonyOS ecosystem (Liu et al., 2024). PLMs/LLMs may not be pre-trained on ArkTS, but have seen massive Java code, including the one to be translated. In this case, a contamination scenario concerning only input portions occurs. Other scenarios are also prevalent in code tasks, such as the unit test generation task. Perhaps both the focal methods and test cases have been pre-trained by PLMs/LLMs, but they are placed in different files on GitHub, leading to their unpaired learning during the pre-training stage of PLMs/LLMs. In this case, how does it affect PLMs/LLMs’ performance on downstream tasks? As such, this work seeks to fill this important gap.
@@
Empirical Study: To explore the contamination effects in a fine-grained manner, we conduct a comprehensive empirical study involving two PLMs and two LLMs on three representative code tasks, including code translation, code generation, and code summarization, encompassing two PLs, i.e., Java and Python. In our analysis, we identify four typical types of data contamination scenarios commonly encountered in code intelligence: (1) input-only contamination, where only the input portion of the test sets appears in the pre-training corpora; (2) output-only contamination, where only the output portion is contaminated; (3) unpaired contamination, where both input and output components are seen in pre-training but not as matching pairs; and (4) paired contamination, where the exact input-output pairs from the test sets are found in the pre-training data of PLMs or LLMs.
@@
As for experiments with PLMs, we adopt a pre-training, fine-tuning, then inference paradigm to simulate their practical usage. To circumvent uncontrolled data leakage, we perform pre-training from scratch (Magar and Schwartz, 2022; Jiang et al., 2024b) and simulate contamination by injecting test data into the pre-training corpus according to different contamination settings. After fine-tuning on downstream tasks, we evaluate model performance on test sets and compare with the uncontaminated counterparts, thereby measuring the contamination effects. As for LLMs, we also follow their practical usage paradigm, i.e., pre-training then inference, to conduct the experiments. We use their publicly available pre-training corpora (Touvron et al., 2023; Li et al., 2023) to extract appropriate samples according to different contamination settings as contaminated testing sets. As for uncontaminated counterparts, we design a series of perturbation rules to modify the contaminated testing sets, making them unseen to LLMs (Mastropaolo et al., 2023) while keeping sample complexities the same. All experiments are repeated five times for statistical analysis, thereby making comparisons and discussing the effects of contamination.
@@
Main Findings: (1) PLMs with pre-training, then fine-tuning and inference paradigm are surprisingly not significantly affected by any contamination settings, yielding on average (-0.07%)–0.2% and (-0.03%)–0.13% fluctuations on BLEU and METEOR over uncontaminated counterparts across three code tasks (p-value0.05). (2) Direct inference or small-scale fine-tuning uncovers the contamination effects on decoder-only PLMs, such as GPT-2. In particular, direct inference induces the most significant overestimation, reaching 54.09% and 39.06% in terms of BLEU and METEOR on average (p-value0.05). (3) Large-scale fine-tuning tends to eliminate the contamination effects of decoder-only PLMs by overwhelmingly learning with underlying knowledge and task alignment specifications. (4) LLMs with pre-training then inference paradigm are only affected by paired contamination, on average obtaining consistent overestimation of 13.41% and 7.39% on both metrics overall (p-value0.05). (5) The above conclusions are applicable for both static- and dynamic-typed PLs, such as Java and Python. This paper makes the following contributions.
@@
(3) We conduct extensive experiments among diverse contamination settings, various PLMs/LLMs on multiple code tasks and programming languages. A series of findings are revealed in this work, shedding light on future research and practice.
2. Experimental Setup
@@
2.1. Research Questions
This paper proposes the following four Research Questions (RQs).
2.1.1. RQ1: Does Input-Only Contamination Lead to Performance Overestimation?
When constructing testing sets, only the input portion is included in the pre-training corpora of PLMs/LLMs, thereby simulating the input-only contamination. This is a prevalent scenario, and the Java-to-ArkTS translation is a typical example, as mentioned in Section 1.
A prior study conducted by Jiang et al. (Jiang et al., 2024b) has claimed that the input-only and output-only contamination leads to performance overestimation for PLMs in Natural Language (NL) tasks, such as SQuAD (Rajpurkar et al., 2016) and MMLU (Hendrycks et al., 2020). However, NL tasks differ significantly from code intelligence tasks, and their experimental designs are flawed. First, the NL tasks carried out in Jiang et al.’s study only involve English text, leading to a relatively high similarity between their inputs and outputs compared with code intelligence tasks that always contain both NLs and code from different PLs. Second, the experiments of Jiang et al. only make comparisons in terms of average performance among only three trials without conducting statistically significant tests, making it hard to eliminate the influence of randomness and determine whether the outperformance of data contamination is consistent and prominent. Third, Jiang et al. only conducted their experiments on one PLM, i.e., GPT-2, without considering different architectures of PLMs and other LLMs. Therefore, the input-only contamination is highly worthwhile to re-evaluate in the code intelligence domain.
2.1.2. RQ2: Does Output-Only Contamination Lead to Performance Overestimation?
@@
2.1.3. RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
@@
This is a kind of complete contamination of testing sets, but the input and output portions are not paired in the pre-training corpora of PLMs/LLMs. Recall the example we mentioned in Section 1, when doing unit test case generation task, both focal methods and test cases may exist in the pre-training corpora of PLMs/LLMs, but their source files are split in the wild. Thus, these code samples will be pre-trained in PLMs/LLMs in a split manner. Considering the prevalence of this setting, unpaired input-output contamination is also meaningful and critical for exploration.
2.1.4. RQ4: Does Paired Contamination Lead to Performance Overestimation?
@@
As another kind of complete contamination, paired input-output contamination represents that both the input and output portions of testing sets have been seen by PLMs/LLMs, and they are concatenated in pairs during the pre-training stage. Many code intelligence tasks are very likely to involve this contamination, such as code generation and completion, as their inputs (i.e., NL or code prefix) are always closely attached in front of corresponding outputs (i.e., code snippets) in the wild, leading to their paired appearance in PLMs/LLMs’ pre-training. Theoretically, paired contamination is the most likely type to cause performance overestimation, and the LLM part has been investigated in (Riddell et al., 2024; Ramos et al., 2024), but PLMs with pre-training, fine-tuning, and inference paradigm have not been explored.
2.2. Models under Test
2.2.1. Pre-trained Language Models (PLMs):
@@
Among this category, an encoder-only model, namely RoBERTa (Liu et al., 2019), and a decoder-only model, namely GPT-2 (Radford et al., 2019), are selected for experiments, as they have covered the most representative architectures of PLMs. We select RoBERTa with 125M parameters and GPT-2 with 150M parameters for experiments according to the load of our computational resources. To facilitate our research, we adopt one typical pre-training task for each PLM during its pre-training stage. RoBERTa is pre-trained with the Masked Language Modeling (MLM) (Devlin et al., 2019) task, while the Causal Language Modeling (CLM) (Radford et al., 2018) task is used for GPT-2. Contamination injections are performed in the pre-training stage. Then, PLMs go through fine-tuning and inference stages to observe whether contamination induces performance overestimation compared with the uncontaminated counterparts.
2.2.2. Large Language Models (LLMs):
@@
As for this category, we select one general LLM, namely LLaMA (Touvron et al., 2023), and one code LLM, namely StarCoder (Li et al., 2023). LLaMA is a family of multilingual LLMs with parameter sizes ranging from 7B to 65B. It is pre-trained on 1.4 trillion tokens of various open-source datasets, where 4.5% of the training base is code extracted from GitHub (17). We select its 33B version for experiments. StarCoder is a Code LLM containing 15.5B parameters, trained on over 1 trillion code tokens. LLMs are experimented with in a pre-training then inference paradigm. Due to the massive scale and substantial training base of LLMs, retraining from scratch is impractical. We curate appropriate samples from their pre-training corpora for the contaminated testing set construction and experiments. The reasons we consider the above two LLMs are two-fold: First, they cover both general and code LLMs, making our exploration comprehensive. Second, they are two of the few open-source LLMs that also release their pre-training corpora, enabling us to construct contaminated testing sets and the uncontaminated counterparts from them.
2.3. Code Tasks and Metrics
2.3.1. Code Translation:
@@
This task involves converting code from one PL to another while keeping the functionality the same, which is critical in scenarios such as migrating legacy systems and adopting modern frameworks. In this work, we focus on code translation from Java to C#, and Python to Java for experiments, denoted as JavaC# and PythonJava for simplicity.
2.3.2. Code Generation:
@@
It is a process of automatically producing source code from a given NL requirement, aiming to reduce manual coding effort while ensuring correctness and efficiency. In this work, we adopt both Java and Python code generation for experiments, which is denoted as NLJava and NLPython for simplicity.
2.3.3. Code Summarization:
@@
This task focuses on generating an NL description for a given code snippet, which is crucial for improving code readability, maintenance, and collaboration among developers. Still, we carry out experiments on both Java and Python programs, which are denoted as JavaNL and PythonNL for simplicity.
2.3.4. Metrics:
@@
To automatically assess the quality of the outputs for the above code tasks, we adopt two widely used evaluation metrics, i.e., BLEU (Papineni et al., 2002) and METEOR (Banerjee and Lavie, 2005). BLEU measures the precision of -gram overlap between the generated output and the reference, focusing on how much of the generated content is correct. In contrast, METEOR incorporates stemming, synonym matching, and alignment-based evaluation, placing more emphasis on recall—i.e., how comprehensively the generated output covers the information in the reference. By combining these two complementary metrics, we provide a more balanced and holistic assessment of generation quality from both precision and recall perspectives.
2.4. Datasets
2.4.1. Dataset preparation for PLMs:
Since all included code tasks are divided into a Java side and a Python side as mentioned in Section 2.3, we use the Java and Python portions of the widely adopted CodeSearchNet dataset (Husain et al., 2019) to pre-train PLMs, respectively, for follow-up exploration accordingly. The Java portion contains over 160,000 Java function snippets, while the Python portion includes approximately 250,000 Python function snippets.
Afterwards, to smoothly carry out experiments on each code task, we carefully select a series of typical datasets. For the code translation task, we adopt the CodeTrans dataset (Lu et al., 2021) for experiments on the Java-to-C# translation, which includes 10,300 training samples, 500 validation samples, and 1,000 test samples. As for the Python-to-Java translation, we use the AVATAR dataset (Ahmad et al., 2023), containing 60,138 training samples, 476 validation samples, and 1,906 testing samples. For the code generation task, we use the CONCODE dataset (Iyer et al., 2018) for NL-to-Java fine-tuning and inference, which contains 100,000 training samples and 2,000 validation and testing samples. For NL-to-Python code generation, we conduct experiments on the Gretel AI dataset (AI, 2024), which consists of 22,500 training samples, 2,500 validation samples, and 2,500 test samples. For the code summarization task, TL-CodeSum dataset (Hu et al., 2018) is selected for experiments on summarizing Java code via NL, consisting of 55,767 training samples and 6,970 validation and testing samples, respectively. In addition, the code-docstring-corpus dataset (Miceli Barone and Sennrich, 2017) is chosen for Python-targeted code summarization, where 109,108 samples are for training, and 2,000 samples are for validation and testing, respectively. Since the above datasets are open-sourced and the corresponding illustrations are detailed in their original papers, we do not list their specific statistics here, thereby saving page limits. To restrict the experimental scales while maintaining the validity of the whole research, we limit the size of the training set to no more than 100,000 samples, the validation set to no more than 500 samples, and the testing set to 1,000 samples at most; when their original volumes exceed the above thresholds, we make corresponding random sampling. Detailed experimental settings and methodologies are presented in Section 2.5.
2.4.2. Dataset preparation for LLMs: